# Is Translation All You Need? A Study on Solving Multilingual Tasks with Large Language Models

### v1

@@ Abstract
Large language models (LLMs) have demonstrated strong multilingual capabilities; yet, they are mostly English-centric due to the imbalanced training corpora. Existing works leverage this phenomenon to improve their multilingual performances on NLP tasks. In this work, we extend the evaluation from NLP tasks to real user queries. We find that even though translation into English can help improve the performance of multilingual NLP tasks for English-centric LLMs, it may not be optimal for all scenarios. For culture-related tasks that need deep language understanding, prompting in the native language proves to be more promising since it can capture the nuances related to culture and language. Therefore, we advocate for more efforts towards the development of strong multilingual LLMs instead of just English-centric LLMs.
@@ A Study on Solving Multilingual Tasks with Large Language Models
[…]
@@ 1 Introduction
Figure 1: Illustration of two kinds of LLMs on different tasks. "English-centric LLMs" refers to LLMs trained mainly in English corpora. "Multilingual LLMs" refers to ideal LLMs that are equally capable in all languages.
@@
To enhance performances in multilingual natural language processing (NLP) tasks with English-centric language models, translating training or test data into English has proven an effective strategy Conneau et al. (2018); Ponti et al. (2020). Recent investigations have expanded this concept by incorporating translation, either implicitly or explicitly, into the intermediate stages of prompting LLMs Huang et al. (2023); Qin et al. (2023); Etxaniz et al. (2023) for multilingual NLP tasks. For example, Shi et al., 2022 demonstrates that translating test questions into English enhances performance on multilingual reasoning tasks, as illustrated in Figure 2(a). Similarly, Huang et al., 2023 and Etxaniz et al., 2023 have shown that prompting LLMs to first translate or comprehend questions in English, then proceed to solve problems step by step, leads to better performance.
@@
Figure 2: Examples illustrating how translation can both improve (top) and degrade (bottom) the performance of LLMs. The example in Chinese is drawn from MGSM and the example in Swahili is drawn from M3Exam, respectively. Translation is beneficial when the questions are language-independent. However, for questions requiring deep cultural knowledge, translation can hinder the ability to answer correctly.
@@
Despite these advancements, the methodologies employed across different studies vary considerably, and the specific impact of translation in improving multilingual task performance has not been rigorously analyzed. Furthermore, the focus of these studies tends to be on specific NLP tasks and English-centric LLMs, with less attention given to real-world scenarios of multilingual user requests. This gap highlights a need for more nuanced research into the effectiveness of translation techniques across a spectrum of multilingual contexts.
In this study, we conduct a thorough investigation into the utility of translation with large language models for a variety of scenarios, including multilingual NLP tasks and real-world multilingual user queries. We compare translating multilingual tasks into English, with an optional step of translating responses back into the original languages against several baselines, such as using native language prompts and cross-lingual prompts Huang et al. (2023). Our experiments also span across a diverse set of LLMs and translation tools to ensure the broad applicability of our findings.
Our results demonstrate that simply translating queries into English can already achieve the best results in multiple task categories. Subsequently, we investigate the utility of translation in real-world scenarios by conducting tests with requests from ShareGPT [^1] https://sharegpt.com/. We discover that translation tends to yield better responses for low-resource languages. However, when working with high-resource languages and advanced LLMs, employing prompts in native languages appears to be the more effective strategy. To further analyze such behaviors, we extend our evaluations to LLMs optimized for languages other than English, focusing specifically on culture-related tasks. We observe that for these models, prompts in the native languages yield superior results by capturing the nuances related to culture and language, as illustrated in Figure 2 (b). In summary, as illustrated in Figure 1, while translation can improve the performance of English-centric LLMs, it may not be optimal for all scenarios. In this work, we highlight the need for robust LLMs that are equally proficient across all languages to effectively address a broader range of cases. Therefore, we advocate for more efforts towards the development of multilingual LLMs.
@@ 2 Translation for NLP Tasks
In this section, we explore various prompting strategies across multiple languages and LLMs, encompassing a wide spectrum of NLP tasks. This allows us to understand how different prompting methods impact the performance of tasks that theoretically do not depend on any specific language.
@@ 2.1 Experiment Setup
Original Question
制作一件袍子需要 2 匹蓝色纤维布料和这个数量一半的白色纤维布料。它一共需要用掉多少匹布料
Native-Basic
{Original Question }
您的最终答案的格式应为:"答案: <阿拉伯数字>".
EN-Basic
{Original Question }
You should format your final answer as "Answer: <Arabic numeral>".
Native-CoT
{Original Question }
让我们一步步思考。
您的最终答案的格式应为:"答案: <阿拉伯数字>".
EN-CoT
{Original Question }
Let’s think step by step in English.
You should format your final answer as "Answer: <Arabic numeral>".
XLT
I want you to act as an arithmetic reasoning expert for Chinese.
Request: {Original Question}
You should retell the request in English.
You should do step-by-step answer to obtain a number answer.
You should step-by-step answer the request.
You should tell me the answer in this format ’Answer :’.
Trans-Google
Crafting a robe requires 2 bolts of blue fiber cloth and half that amount of white fiber cloth. How many pieces of fabric will it use in total?
Let’s think step by step.
You should format your final answer as "Answer: <Arabic numeral>".
Trans-NLLB
To make a robe, two pieces of blue fiber and half of that amount of white fiber are needed. How many pieces of fabric does it take to make?
Let’s think step by step.
You should format your final answer as "Answer: <Arabic numeral>".
We perform assessments on six benchmarks on reasoning, understanding, and generation tasks that can encapsulate various abilities of LLMs. We follow the task and language selection from Huang et al. (2023) to consider a wide array of 24 diverse languages, including German (de), Russian (ru), French (fr), Chinese Simplified (zh), Spanish (es), Japanese (ja), Italian (it), Vietnamese (vi), Turkish (tr), Indonesian (id), Swahili (sw), Arabic (ar), Korean (ko), Greek (el), Thai (th), Bulgarian (bg), Hindi (hi), Estonian (et), Bengali (bn), Tamil (ta), Urdu (ur), Telugu (te), Haitian Creole (ht), and Southern Quechua (qu). We categorize languages larger than 1% frequency in Common Crawl [^2] https://commoncrawl.github.io/cc-crawl-statistics/plots/languages as high-resource languages (i.e., de, ru, fr, zh, es, ja, it and vi), and the rest as low-resource languages. We exclude English since we want to evaluate the efficient prompting strategy for non-English tasks. As our prompting strategies involve methods with translators, we do not test on the FLORES-200 dataset.
@@ 2.1.1 Tasks
We evaluate the models and prompting strategies on the following tasks: 1) the MGSM benchmark Shi et al. (2022), which tests models on grade school math problems. 2) the XCOPA benchmark Ponti et al. (2020), which presents a premise with two choices to determine causality or effect; 3) the XNLI benchmark Conneau et al. (2018), which assesses Natural Language Inference by judging the relationship between a premise and a hypothesis; 4) the PAWS-X benchmark (Yang et al., 2019) involves identifying paraphrases in two sentences; 5) the MKQA dataset (Longpre et al., 2021), which focuses on open-domain questions; 6) the XL-Sum benchmark Hasan et al. (2021), which involves summarizing news articles. We choose a subset of 9 languages for MKQA and 5 languages for XL-Sum. For evaluation metrics across our study, we employ the token overlap F1 score specifically for the MKQA dataset, the ROUGE-1 score for assessing XL-Sum, and accuracy as the standard metric for all other benchmarks. More details of the benchmarks can be found in A.1.
For each task, we sample up to 500 examples from the test set for each language. For MGSM, since there are only 250 examples for each language, the entire test set will be used. For generation tasks like MKQA and XL-Sum, we will translate the answer back to the original language if the prompting strategy involves a translator.
@@ 2.1.2 Models
We conduct experiments with six LLMs, consisting of one closed-source language model and five open-source language models:
  • ChatGPT [^3] https://platform.openai.com/docs/models/gpt-3-5, which is the most capable and cost-effective model in the GPT-3.5 family and it has been optimized for chat. We chose the latest version (gpt-3.5-turbo-1106) when starting the experiment.
  • Llama-2-13B-chat and Llama-2-70B-chat, which are the chat models in Llama-2 family Touvron et al. (2023). Due to computational resource limitations, we use the AWQ Lin et al. (2023) version for Llama-2-70B-chat.
  • bloomz-7b1, which is a model fine-tuned with multiple tasks, including multilingual tasks Muennighoff et al. (2023).
@@ 2.1.3 Prompting Strategies
Considering the language of instruction, whether to apply a chain-of-thought reasoning approach and the availability of translation tools, we assess the following experimental strategies in the experiment. We utilize a zero-shot approach across all strategies, given that the models we selected have been fine-tuned to follow instructions.
@@
XLT Huang et al. (2023) is a state-of-the-art prompting method to handle multilingual NLP tasks. It prompts LLMs to translate the question into English and solve the problem step-by-step in English, leveraging the English-centric bias of LLMs.
@@
In addition to the prompting strategies, an output constraint is also included in the template to facilitate answer extraction. The settings are demonstrated in Table 1. In the case of open-source models, where the output format may deviate from the instructions, we utilize "Therefore, the answer <constraint> is" in appropriate languages in the second round to retrieve the ultimate answer.
2.2 Results and Discussions
Model Prompt type MGSM XCOPA XNLI PAWS-X MKQA XL-Sum AVG
ChatGPT Native-Basic 34.4 72.3 52.2 49.7 35.4 31.9 46.0
EN-Basic 41.1 76.1 63.0 62.0 36.5 32.7 51.9
Native-CoT 49.9 72.4 50.6 50.3 35.4 31.1 48.3
EN-CoT 61.1 78.6 56.4 61.6 42.9 29.7 55.1
XLT 62.3 79.3 59.2 59.4 37.6 23.4 53.5
Trans-Google 73.9 84.5 60.5 67.2 43.8 30.3 60.0
Trans-NLLB 61.0 79.7 59.2 67.5 37.2 28.3 55.5
bloomz-7b1 Native-Basic 1.3 21.4 8.3 - 7.8 - 9.7
EN-Basic 2.0 56.5 43.9 - 10.6 - 28.3
Native-CoT 1.2 20.9 8.2 - 6.5 - 9.2
EN-CoT 1.7 53.9 35.9 - 9.3 - 25.2
XLT 1.7 50.5 35.4 - 8.0 - 23.9
Trans-Google 2.7 63.7 44.3 - 17.2 - 32.0
Trans-NLLB 2.4 61.8 43.8 - 14.7 - 30.7
Mistral-7B-Instruct Native-Basic 11.2 54.2 42.8 44.6 7.8 25.9 31.1
EN-Basic 23.8 34.9 50.2 46.9 7.0 21.6 30.7
Native-CoT 17.0 53.8 43.4 44.3 7.8 24.6 31.8
EN-CoT 27.6 40.8 50.0 46.6 11.5 18.7 32.5
XLT 31.8 61.5 46.0 47.8 9.6 17.6 35.7
Trans-Google 41.3 59.2 55.0 51.5 17.0 32.8 42.8
Trans-NLLB 31.7 54.4 53.0 52.4 15.5 32.2 39.9
Llama-2-13b-Chat Native-Basic 15.6 50.1 36.4 54.0 18.3 27.5 33.6
EN-Basic 19.0 54.3 43.4 59.1 20.2 31.1 37.9
Native-CoT 18.1 50.9 35.7 54.8 15.7 27.2 33.7
EN-CoT 19.9 54.5 43.7 57.6 19.8 30.2 37.6
XLT 22.3 56.0 51.4 55.7 19.0 21.2 37.6
Trans-Google 39.1 71.9 46.1 58.4 33.8 35.7 47.5
Trans-NLLB 31.8 68.2 45.4 57.8 28.4 34.2 44.3
Llama-2-70B-Chat Native-Basic 23.6 51.5 39.0 52.8 24.8 29.4 36.2
EN-Basic 28.6 55.3 46.5 60.4 24.7 31.2 40.1
Native-CoT 23.5 50.5 37.9 54.9 21.9 25.8 34.5
EN-CoT 30.2 61.2 45.9 64.9 31.1 30.2 42.3
XLT 32.8 58.7 52.2 55.7 26.6 17.7 38.4
Trans-Google 53.3 80.9 54.0 68.5 39.7 33.8 51.8
Trans-NLLB 43.8 77.1 52.2 69.2 19.4 33.1 49.1
Table 2: Average scores across different languages (excluding English) for the six benchmarks in zero-shot setting. The results of PAWS-X and XL-Sum for bloomz-7b1 are not considered since it was already pre-trained on these tasks. The best result for each model is in bold.
@@ 2.2 Results and Discussions
The main results are shown in Table 2. We also calculate the average performance for high-resource and low-resource languages separately for each task in Table 4 in Appendix A.2. We notice that Trans-Google, despite simple, surprisingly demonstrates the highest overall performance across various models and tasks. While it may not always achieve top performance, it consistently delivers commendable results for both high and low-resource languages. Besides this, we can have the following observations: 1) Trans-NLLB ranks second on average, trailing behind Trans-Google. The discrepancy between the two methods lies solely in the quality of translation, indicating that the NLLB model’s translation proficiency is inferior. Nonetheless, its open-source nature renders it a potentially favourable choice under specific circumstances. 2) Utilizing English instructions generally enhances performance across various tasks, regardless of the integration of chain-of-thought. This finding aligns with those reported by Lai et al., 2023. 3) chain-of-thought is quite helpful for strong LLMs like ChatGPT and reasoning tasks like MGSM. For weaker models and tasks that can be answered directly, the basic prompt may be a better option. More results and discussions are in Appendix A.2.

### v2

@@ Abstract
Large language models (LLMs) have demonstrated multilingual capabilities; yet, they are mostly English-centric due to the imbalanced training corpora. Existing works leverage this phenomenon to improve their multilingual performances through translation, primarily on natural language processing (NLP) tasks. This work extends the evaluation from NLP tasks to real user queries and from English-centric LLMs to non-English-centric LLMs. While translation into English can help improve the performance of multilingual NLP tasks for English-centric LLMs, it may not be optimal for all scenarios. For culture-related tasks that need deep language understanding, prompting in the native language tends to be more promising as it better captures the nuances of culture and language. Our experiments reveal varied behaviors among different LLMs and tasks in the multilingual context. Therefore, we advocate for more comprehensive multilingual evaluation and more efforts toward developing multilingual LLMs beyond English-centric ones.
@@ A Study on Solving Multilingual Tasks with Large Language Models
Chaoqun Liuthanks: Chaoqun Liu is under the Joint PhD Program between DAMO Academy and Nanyang Technological University.   12  Wenxuan Zhangthanks: Wenxuan Zhang is the corresponding author.  23  Yiran Zhao24  Anh Tuan Luu1  Lidong Bing23 1Nanyang Technological University, Singapore; 2DAMO Academy, Alibaba Group, Singapore 3Hupan Lab, 310023, Hangzhou, China; 4National University of Singapore {chaoqun.liu,saike.zwx,zhaoyiran.zyr,l.bing}@alibaba-inc.com anhtuan.luu@ntu.edu.sg
@@ 1 Introduction
Figure 1: Illustration of two types of LLMs on tasks with varying language dependencies. "English-centric LLMs" refers to LLMs trained mainly in English corpora. "Multilingual LLMs" refers to ideal LLMs equally capable in all languages.
@@
To enhance performances in multilingual natural language processing (NLP) tasks with English-centric language models, translating training or test data into English has proven an effective strategy Conneau et al. (2018); Ponti et al. (2020); Artetxe et al. (2023); Moghe et al. (2023). Recent investigations have expanded this idea by incorporating translation, either implicitly or explicitly, into the intermediate stages of prompting LLMs Huang et al. (2023); Qin et al. (2023b); Etxaniz et al. (2023) for multilingual NLP tasks. For example, Shi et al., 2022 demonstrates that translating test questions into English enhances performance on multilingual reasoning tasks, as illustrated in Figure 2(a). Similarly, Huang et al., 2023 and Etxaniz et al., 2023 have shown that prompting LLMs to first translate or comprehend questions in English, then solve them step by step, improves performance.
@@
Figure 2: Examples illustrating how translation can both improve (a) and degrade (b) the performance of LLMs. The Chinese example is from MGSM Shi et al. (2022) and the Swahili example is from M3Exam Zhang et al. (2023a). Translation is beneficial when the questions are semantically equivalent across languages. However, for questions that demand deep cultural knowledge, translation can hinder the ability to answer accurately.
@@
Despite these advancements, methodologies in various studies differ significantly, and the impact of translation on multilingual task performance remains underexplored. Furthermore, these studies focus on specific NLP tasks and English-centric LLMs, with less attention given to real-world user requests in various languages. This gap highlights a need for more nuanced research into the effectiveness of translation techniques across multilingual contexts. As shown in Figure 1, we hypothesize that English-centric LLMs generally perform better with English translations of prompts, while "Multilingual LLMs" excel with native prompts, particularly for tasks highly dependent on language.
In this study, we perform an in-depth analysis of the utility of translation with large language models for various scenarios. Firstly, we compare translating multilingual tasks into English, with an optional step of translating responses back into the original languages (i.e., the “translate-test” method), against several baselines on multilingual NLP tasks. Secondly, we extend the evaluation to real user queries, which are more likely to contain knowledge related to culture and language. Thirdly, We broaden the scope of LLM evaluations to include non-English-centric models to explore how they differ in behavior from English-centric LLMs.
Our results demonstrate that simply translating queries into English can already achieve the best results in multiple NLP task categories. For real user queries, the effect of translation depends on the languages and the LLMs. When working with advanced LLMs and certain languages, employing prompts in native languages appears to be the more effective strategy. In addition, the non-English-centric LLMs also behave differently from English-centric LLMs, where prompts in the native languages yield superior results by capturing the nuances related to culture and language.
The main contributions of this work can be summarized as follows:
  • We compare various multilingual prompting strategies in NLP tasks, finding that translation remains a strong baseline even for LLMs. Our study also identifies factors that affect performance in multilingual tasks.
  • We broaden our multilingual evaluation to encompass actual user queries and LLMs that are not primarily centered on English, thereby addressing the limitations of previous works.
  • Our study reveals the shortcomings of current multilingual evaluations, underscoring the need for more comprehensive assessments that broaden the range of benchmarks and diversify the types of LLMs considered.
@@ 2 Translation for NLP Tasks
This section explores various prompting strategies across multiple languages and LLMs, covering a wide range of NLP tasks. This helps us understand how different prompting methods and other factors affect task performance.
@@ 2.1.1 Tasks
We conduct assessments on six benchmarks covering reasoning, understanding, and generation tasks that encapsulate various abilities of LLMs: MGSM Shi et al. (2022), XCOPA Ponti et al. (2020), XNLI Conneau et al. (2018), PAWS-X (Yang et al., 2019), MKQA (Longpre et al., 2021) and XL-Sum Hasan et al. (2021). Following Huang et al., 2023, we choose a subset of 9 languages for MKQA and 5 languages for XL-Sum. For evaluation metrics across our study, we employ the token overlap F1 score specifically for the MKQA dataset, the ROUGE-1 score for assessing XL-Sum, and accuracy as the standard metric for all other benchmarks. More details of the benchmarks can be found in Appendix A.1.
These tasks cover a wide array of 24 diverse languages, including German (de), Russian (ru), French (fr), Chinese Simplified (zh), Spanish (es), Japanese (ja), Italian (it), Vietnamese (vi), Turkish (tr), Indonesian (id), Swahili (sw), Arabic (ar), Korean (ko), Greek (el), Thai (th), Bulgarian (bg), Hindi (hi), Estonian (et), Bengali (bn), Tamil (ta), Urdu (ur), Telugu (te), Haitian Creole (ht), and Southern Quechua (qu). We categorize languages larger than 1% frequency in Common Crawl [^1] https://commoncrawl.github.io/cc-crawl-statistics/plots/languages as high-resource languages (i.e., de, ru, fr, zh, es, ja, it and vi), and the rest as low-resource languages. We exclude English since we want to evaluate the efficient prompting strategy for non-English tasks.
For each task, we sample 500 examples from the test set per language or use the entire test set if there are fewer than 500 examples. For generation tasks like MKQA and XL-Sum, answers will be translated back to the original language if the prompting strategy uses a translator.
Model Prompt type MGSM XCOPA XNLI PAWS-X MKQA XL-Sum AVG
high low high low high low high low high low high low high low
XLT 70.4 50.1 89.3 76.8 60.6 58.1 59.7 58.2 37.7 37.5 22.8 26.1 56.7 51.1
XLT 49.0 8.4 76.4 54.7 57.3 48.4 56.6 51.6 26.5 26.7 19.3 11.5 47.5 33.6
Table 1: Average scores of the high-resource languages and low-resource languages for the six benchmarks in zero-shot setting. The best result for each model is in bold.
@@ 2.1.2 Models
We mainly conduct experiments on the following two LLMs, consisting of one closed-source language model and one open-source language model:
ChatGPT
This is the most capable and cost-effective model in the GPT-3.5222https://platform.openai.com/docs/models/gpt-3-5 family optimized for chat. We chose the latest version (gpt-3.5-turbo-1106) when starting the experiment.
Llama-2-70B-Chat
This is the largest chat models in Llama-2 family Touvron et al. (2023). Due to computational resource limitations, we use the AWQ Lin et al. (2023) version for evaluation.
We also conducted experiments on some other models, including Mistral-7B-Instruct (v0.2) Jiang et al. (2023), Llama-2-13B-chat Touvron et al. (2023) and bloomz-7b1 Muennighoff et al. (2023). More details are shown in Appendix A.1.
@@ 2.1.3 Prompting Strategies
Considering the language of instruction, the use of chain-of-thought reasoning, and translation tools, we assess the following experimental strategies. We employ a zero-shot approach for all strategies, given that our selected models are fine-tuned to follow instructions.
@@
XLT Huang et al. (2023) is a state-of-the-art prompting method to handle multilingual NLP tasks. It prompts LLMs to translate the question into English and solve the problem step-by-step in English.
@@
The examples for each strategy are shown in Table A.1.4 and the templates for EN-Basic are shown in Table 5 in the Appendix. In addition to the prompting strategies, an output constraint is also included in the template to facilitate answer extraction. When the output format may deviate from the instructions, we utilize "Therefore, the answer <constraint> is" in appropriate languages in the second round to retrieve the ultimate answer.
2.2 Main Results
@@
The main results are shown in Table 1. We notice that Trans-Google, despite simple, demonstrates the highest overall performance across various models and tasks. While it may not always achieve top performance, it consistently delivers commendable results for both high and low-resource languages. Besides this, we can have the following observations: 1) Utilizing English instructions generally enhances performance across various tasks, regardless of the integration of chain-of-thought. This finding aligns with those reported by Lai et al., 2023. 2) chain-of-thought is quite helpful for strong LLMs like ChatGPT and reasoning tasks like MGSM. For weaker models and tasks that can be answered directly, the basic prompt may be a better option. 3) On average, EN-CoT underperforms compared to Trans-Google for both high and low-resource languages. While EN-CoT surpasses Trans-NLLB in high-resource languages, it falls short in low-resource ones. We hypothesize that this discrepancy arises because LLMs excel in high-resource languages but need external translation systems to handle low-resource languages effectively. More results and discussions are in Appendix A.1.4.
2.3 Analysis and Discussions
To investigate the impact of different factors on performance across various languages, we conduct a series of experiments and analyses using the MGSM benchmark.
Is there a relationship between task performance and translation quality?
In addition to external translation systems, we can use LLMs to translate the questions. Although XLT includes translation, it is integrated into the solutions. Therefore, we examine the self-translate approach Etxaniz et al. (2023), translating in a zero-shot manner with the prompt template shown in Appendix A.1.3. Then we prompt LLMs with the translated question the same as Trans-Google and Trans-NLLB. The results are shown in Table 8 in the Appendix.
We use the English subset of MGSM as the reference translation and evaluate translation quality using the SacreBLEU score Papineni et al. (2002); Post (2018). The results, shown in Figure 3, indicate that Google Translate achieves the highest quality for all languages except Japanese. Translations by ChatGPT (Trans-ChatGPT) and Llama-2-70B-Chat (Trans-Llama) outperform Trans-NLLB for high-resource languages but not for some low-resource languages.
Figure 3: BLEU scores for translating MGSM questions with different translation systems.
To analyze the impact of translation quality on final performance, we plot the correlation between accuracy scores and BLEU scores for each language in Figure 4. The results show that higher translation quality (BLEU scores) generally leads to better task performance, highlighting the importance of an effective translation system.
Figure 4: Corrections between BLEU scores of translation and MGSM accuracy for the three prompting techniques: Trans-Google, Trans-NLLB and self-translate. Each dot in the figure represents the performance of one model on one language.
Does language distance between English and target language affect the performances?
Table 1 shows that the LLMs perform better for high-resource languages than low-resource languages on average. We hypothesize that language distance, besides language frequency, is crucial for English-centric LLMs. To verify this, we calculate the correlation between MGSM accuracy and the language distances between the target languages and English. Following Philippy et al., 2023, We examine five types of distances, including the syntactic (SYN), geographic (GEO), inventory (INV), genetic (GEN), and phonological (PHON) distances extracted using lang2vec Littell et al. (2017). As shown in Table 2, MGSM accuracy significantly correlates with syntactic distance but not with other types of distances. The negative values indicate that languages with a larger syntactic distance from English tend to perform worse.
Prompt type SYN GEO INV GEN PHON
ChatGPT
Native-Basic -0.786* -0.336 0.323 -0.403 -0.044
EN-Basic -0.820* -0.160 0.527 -0.299 0.020
Native-CoT -0.795* -0.184 0.479 -0.313 0.045
EN-CoT -0.841* -0.286 0.339 -0.436 -0.034
XLT -0.787* -0.113 0.445 -0.284 0.117
Llama-2-70B-Chat
Native-Basic -0.688* -0.369 0.250 -0.323 -0.044
EN-Basic -0.782* -0.512 0.134 -0.513 -0.226
Native-CoT -0.706* -0.403 0.231 -0.475 -0.105
EN-CoT -0.737* -0.510 0.206 -0.445 -0.219
XLT -0.697* -0.432 0.266 -0.423 -0.153
Table 2: Pearson correlation coefficient between MGSM accuracy and five language distances between English and that language. (*p < 0.05, two-tailed)